why is x=-1 the solution of (x-2)*(3x) + (2x-1)*(x+4) = (5x-1)*(x+1)?
Found 2 solutions by ewatrrr, Theo:
Answer by ewatrrr(24785)


Hi
(x-2)(3x) + (2x-1) (x+4) = (5x-1) (x+1)?
Expanding in order to combine Like terms:
(x-2)(3x) + (2x-1)(x+4) = (5x-1)(x+1)
3x^2-6x + 2x^2+7x-4 = 5x^2+4x-1
x -4 = 4x -1 |subtracting 5x^2 from both sides of the Eq
-3 = 3x
-1 = x

Answer by Theo(13342)

when you substitute -1 for x, you get the following:

original equation:

(x-2)*(3x) + (2x-1)*(x+4) = (5x-1)*(x+1)

when you replace x with -1:

(x-2)*(3x) becomes -3*-3 which equals 9
(2x-1)*(x+4) becomes -3*3 which equals -9
(5x-1)*(x+1) becomes -6*0 which equals 0

the equation becomes:
9 - 9 = 0 which becomes 0 = 0 which is true.

they are equal because when you substitute (-1) for x, the equation becomes 0 = 0 which is true.